OpenStudy (anonymous):

Simplify the following expression completely: (-2 y^19)^4

OpenStudy (moonlitfate):

This is just power of powers (I hope I got that right) and the distributive property. :)

OpenStudy (moonlitfate):

\[(-2y ^{19})^4\]

OpenStudy (moonlitfate):

\[=-2^4 *y ^{19*4}\]

OpenStudy (moonlitfate):

@noemi -- Does that help a bit?

OpenStudy (anonymous):

no i don't understand

OpenStudy (moonlitfate):

Okay, well, what exactly don't you understand?

OpenStudy (anonymous):

are you sopose to times the 19 and 4 and put it \[-2^4*y76\]

OpenStudy (anonymous):

or is it going to be \[-2y^80\]

OpenStudy (moonlitfate):

When you have an exponent raised to another exponent, the two exponents are multiplied, if that makes sense.

OpenStudy (moonlitfate):

\[(y ^{19})^4 = y ^{19(4)} = y ^{76}\]

OpenStudy (moonlitfate):

And since -2 is in front of the y, you are also going to raise it to the 4th power.

OpenStudy (anonymous):

so you would multiply 4*76 =304 so my problem would be\[-2y^3^0^4\]

OpenStudy (anonymous):

\[-2y^304\]

OpenStudy (moonlitfate):

\[(-2y ^{19})^4 = -2^4*y ^{19*4} = 16y ^{76}\]

OpenStudy (moonlitfate):

The 76 doesn't get multiplied by anything.

OpenStudy (anonymous):

oh make sense

OpenStudy (moonlitfate):

Try another one. :) I'll just make up something.

OpenStudy (moonlitfate):

\[(-5x^7)^6\]

OpenStudy (anonymous):

its\[-78125x^42\]

OpenStudy (moonlitfate):

Your x exponent is right; the constant isn't. o:

OpenStudy (moonlitfate):

\[-5^6 = 15625\]

OpenStudy (anonymous):

oh ur right i multiplied the wrong one

OpenStudy (anonymous):

\[(4z^19)^4 = 256z^76\]